40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Message ID: 107     Entry time: Tue Apr 13 18:23:41 2010
Author: Tara Chalermsongsak 
Type: Electronics 
Category:  
Subject: TF of PDH Box 

I measured TF of PDH box, D0901351, (The one we have was modified). This box sends the signal to VCO.

 

SR785 measures at low frequency ( 1 Hz to 100kHz)

4935A measures at high frequency (10Hz to 1Mhz)

The integrator switch of the PDH box is turned off. This will be calculate later. The gain is set at 10.

The magnitude as mesured by 4935A is corrected for impedence match by x1.2.( 4935A and the PDH box have 50 ohm impedence for both inputs and outputs.)

This data will be used for control loop model later.

Blue, data from sr785

Green, Data from 4395A, I didnot use the power splitter to split the signal from source.

Red, Data from 4395A, with power splitter to divide power from source. (The power has to be increased to -30dB)

The first plot is magnitude of the TF, the second plot is phase shift, as usual Bode plot.

Attachment 1: tf_PDH_04_09.png  17 kB  | Hide | Hide all
tf_PDH_04_09.png
Attachment 2: tf_PDH_04_09_phi.png  18 kB  | Hide | Hide all
tf_PDH_04_09_phi.png
ELOG V3.1.3-